Circular No. 8480

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005ae IN ESO 209-9 Further to IAUC 8353, R. Martin reports the discovery of an apparent

supernova (mag 17.5) on images taken on Feb. 1.598 and 2.603 UT. SN 2005ae is located at α = 7^{h}58^{m}15^{s}.84, δ = −49^{o}51′19′′.9 (equinox 2000.0), which is 13′′.5 east and 13′′.9 south of the center of ESO 209-9 and in an area where there is substantial background luminosity from the galaxy. Nothing was visible at this location on an image taken on Jan. 31.661 (limiting magnitude g ≈ 18.0). Martin reports the following magnitudes obtained of the new object on Feb. 8.520: V = 15.9, R = 15.4, I = 14.7.

SUPERNOVA 2005ab IN NGC 4617 S. Benetti writes: “F. Di Mille, Universitá di Padova, on behalf of

the ‘RTN Winter School on Supernovae’ in Asiago, reports that a noisy spectrum (range 390–815 nm, resolution 2.3 nm) of SN 2005ab (cf. IAUC 8478), taken on Feb. 9.00 UT with the Asiago 1.22-m telescope (+ Boller & Chivens spectrograph), is that of a type-II supernova shortly after explosion. The spectrum consists of a blue continuum overimposed by a relatively broad (5400 km/s) Hα emission line.”

COMET C/2004 Q2 (MACHHOLZ) J. H. Sastri and R. Vasundhara, Indian Institute of Astrophysics, Ban-

galore, reports that R-band CCD images of comet C/2004 Q2 were obtained in January by K. Kuppuswamy and C. Velu with the 1.02-m f/13 telescope at the Vainu Bappu Observatory, Kavalur, which reveal dust fans via the spatial filter of Larson and Sekanina (1984, A.J. 89, 571) with the following lengths (± 15′′) and position angles (± 10^{o}) for each of the three fans: Jan. 2.6625 UT, fan 1, 150′′ in p.a. 291^{o}; fan 2, 150′′ in p.a. 252^{o}; fan 3, 60′′ in p.a. 216^{o}. Jan. 15.6344, fan 1, 150′′ in p.a. 40^{o}; fan 2, 150′′ in p.a. 252^{o}; fan 3, 60′′ in p.a. 216^{o}. The dust features were modeled after Vasundhara (2002, A.Ap. 382, 342), indicating that the latitude ranges of the active regions on the nucleus that produce the fans are as follows (direction of fitted north rotational pole α = 190^{o} ± 10^{o}, δ = +50^{o} ± 10^{o}, equinox 2000.0): fan 1, −15^{o} to 0^{o}; fan 2, −50^{o} to −35^{o}; fan 3, −78^{o} to −70^{o}. Assuming silicate grains of size 0.1–30 μm and Fulle’s (1987, A.Ap. 171, 327) relation between grain velocity and the forces on each grain, a rotation period of 0.38 ± 0.08 day is estimated for the comet’s nucleus.

Naked-eye total-magnitude estimates by J. Gonzalez, Asturias, Spain: Jan. 9.81 UT, 3.4; 17.07, 3.7; 22.80, 4.1; 31.95, 4.3; Feb. 7.81, 4.8.

2005 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green